OpenStudy (ny,ny):

HELP!!!!!! Solve each system using substitution. Check your solution. m=4n+11 -6n+8m=36

OpenStudy (anonymous):

place m=4n+11 into -6n + 8m = 36

OpenStudy (anonymous):

Since you already know what m equals, substitute it into the other equation like so -6n + 8(4n+11) = 3

OpenStudy (anonymous):

Then solve for n using algebra, then plug n into either of the original two equations to solve for m

OpenStudy (ny,ny):

-6n+32n+88=36

OpenStudy (anonymous):

right

OpenStudy (anonymous):

now you have this it is just like the other one we did

OpenStudy (ny,ny):

26n=-52

OpenStudy (anonymous):

hold on checking

OpenStudy (anonymous):

right now divide

OpenStudy (ny,ny):

n=-2?

OpenStudy (anonymous):

right now place -2 back into one of the original equation and you have y

OpenStudy (anonymous):

oops m

OpenStudy (ny,ny):

m=3?

OpenStudy (anonymous):

right

OpenStudy (ny,ny):

cool. thanks again moser :D

OpenStudy (anonymous):

welcome you have it just trust yourself

OpenStudy (ny,ny):

ok. I'll try :D